OpenStudy (anonymous):

squareroot of 6 * square root of 2

OpenStudy (jadeishere):

well, find the square root of 6, then multiply it by the square root of 2 rounded its about 3.5

OpenStudy (anonymous):

The answerrs do not come into deciamls i did it that way also, they have 2 square of 3, square of 12, 3 square root of 2, or not possible

OpenStudy (jadeishere):

well then it'd be 2 sqrt of 3

OpenStudy (jadeishere):

the exact form would be \[^{2\sqrt{3}}\] the decimal form would be 3.46

OpenStudy (anonymous):

so it says \[\sqrt{6}*\sqrt{2}\] i wouldnt just mulitply and them and get \[\sqrt{12}\]

OpenStudy (jadeishere):

Nononono

OpenStudy (anonymous):

oh thanks, thanks A LOT ! I struggle with math badly

OpenStudy (jadeishere):

Lol anytime :)